OpenStudy (anonymous):

An isosceles triangle whose base measures 3 and its legs measures 5 find the perimeter pls help

OpenStudy (anonymous):

what is the answer? 6.5?

ganeshie8 (ganeshie8):

perimeter is length around the shape. for a triangle , perimeter = side1 + side2 + side3 = 3 + 5 + 5 = 13

OpenStudy (anonymous):

thanks ^^

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

perimeter = side1 + side2 + side3 = 3 + 5 + 5 = 13

ganeshie8 (ganeshie8):

@Faye281 here the question has some ambiguity. 1) it doesnt say if the SUM of two legs is 5. or 2) each leg is 5 if it is 2nd one. then the perimeter is simply : 3 + 5 = 7

ganeshie8 (ganeshie8):